OpenStudy (anonymous):

x + (10/x-2) = x^2+3x/x-2 solve for x...

OpenStudy (fibonaccichick666):

Let's make sure we have the right problem first: is this it?\[x + \frac{10}{x-2} = \frac{x^2+3x}{x-2}\]

OpenStudy (anonymous):

yes!

OpenStudy (fibonaccichick666):

alrighty so do you have any ideas on how to start?

OpenStudy (anonymous):

nope..

OpenStudy (fibonaccichick666):

ok, how about we start by making the left sdie one big fraction. Can we do that?

OpenStudy (fibonaccichick666):

Can you* do that for me please?

OpenStudy (anonymous):

do i make it say x+10/x^2-2?

OpenStudy (fibonaccichick666):

uhm not quite, tell me what you did there

OpenStudy (anonymous):

i added x to the top and bottom of the fraction. what was i suppose to do?

OpenStudy (fibonaccichick666):

Rules for fractions say this: \[\frac{a}{b}+\frac{c}{d}=\frac{ad+cb}{bd}\]

OpenStudy (anonymous):

ok so i cross multiply then multiply the denominator?

OpenStudy (fibonaccichick666):

no, you multiply each by 1/1 to get a common denominator

OpenStudy (fibonaccichick666):

d/d and b/b =1 hence you can do that

OpenStudy (anonymous):

ok im following along

OpenStudy (fibonaccichick666):

so tell me what you get

OpenStudy (anonymous):

lets swap answers.. lets see what you got first lol

OpenStudy (fibonaccichick666):

nope

OpenStudy (fibonaccichick666):

that's not how I help